Desalo Logo : Website Design & Internet Consultancy

Take pride in your web presence.
Take pride in your web presence

technical support enquiries

For technical support, please use this address:

Service Status

current status: no known issues

  • call 01896 820255 during normal UK working hours,

    or (at any time):
  • SMS +447850176953
  • email

For non-emergency technical support, or for sales enquiries, please:

  • call 01896 820161 or 01896 820255 during normal UK working hours,

  • email


last known service issues:

resolved: hardware failure and resultant server move affecting some sites, but not email (since 20:00, 20/12/2012)

last updated: 21/12/2012, 23:59 (UK local time)


23:59 on 21/12/2012: All back to normal. I've been monitoring things very closely again for the last couple of hours, and all of the sites that were on the affected server last night have been moved and fully restored, as far as I can tell (a couple required some re-uploading from local copies as they'd had updates that were more recent than their backup), and even the sites that remained problematic today, from time to time, are now loading normally again. Response and page-load times also seem to be back to normal for all.

If you do happen to notice any recently-updated page or recently-uploaded image or other file that's missing on your site, please try uploading again from a local copy, if you have one.

If you don't have a local copy (perhaps because I usually carry out updates for you), please let me know on [email protected]. Many thanks.

22:50 on 21/12/2012: It looks like everything is at last pretty much back to normal for all sites now, and most seem to have been fine for a few hours already, according to our monitoring software. If you happen to find that yours is missing any recent updates or uploads, or if you're still experiencing slower load times or even timeouts/lack of availability, please email me to let me know on [email protected] or give me a call on 01896 820161.

Thanks again for all your patience and understanding, everyone.

13:25 on 21/12/2012: As sites are being restored from backup, any recent changes you may have made may not be reflected when your site is restored. If this is the case, please upload your pages again from your local copy, or ask me (Chris) to if I usually maintain your site and do updates (rather than you editing it yourself).

If unsure, please feel free to email me on [email protected] or give me a call on 01896 820161 so we can look at what's best to do to get your own site back to normal as quickly as we can.

10:45 on 21/12/2012: Nearly all sites are now restored from backups. Availability of some restored sites is still fluctuating due to timeouts resulting from the increased server load (due to restoring and related updates still going on). Sorry if your site is one of those affected, and once again, many thanks for your patience and understanding.

If your site is persistently not accepting connections or is down every time you try to visit it, please let me know via [email protected] and I'll get engineers to take a look into things for that specific site.

10:20 on 21/12/2012: We're still watching progress. Availability of some restored sites is fluctuating due to timeouts. This will eventually resolve itself once the server is less busy, but behaviour could well be slow or erratic over the next few hours.

01:50 on 21/12/2012: We're still watching progress and yet more sites are back up, but still not all.

The server team are saying that the move is done and the restore process is stable, but that the high server load while data is being transferred is likely to cause timeouts and slow page loads while the restore is still going on at a pace.

If you happen to notice that your site is persistently not accepting connections or is down every time you try to visit it, please let me know via [email protected] and I'll get engineers to take a look into things for specific sites.

Thanks again for your patience.

On the up side, lots of sites appear to be performing normally.

22:33 on 20/12/2012: More sites are back up. Things are clearly progressing, though I can see that a number of sites are still down, unfortunately.

Will report back here as soon as we hear more from the server fixing team. They seem pretty busy at the moment.

21:58 on 20/12/2012: Some of the affected sites are back up. Ones that are still down may be reported as unreachable, or might display any of a number of standard error messages including:

Internal Server Error

The server encountered an internal error or misconfiguration and was unable to complete your request.

Please contact the server administrator, [email protected][whatever your domain name is] and inform them of the time the error occurred, and anything you might have done that may have caused the error.

More information about this error may be available in the server error log.

Additionally, a 500 Internal Server Error error was encountered while trying to use an ErrorDocument to handle the request.


Gateway Time-out

The gateway did not receive a timely response from the upstream server or application.


No data received

Unable to load the webpage because the server sent no data.

Here are some suggestions:
Reload this webpage later.

Error 324 (net::ERR_EMPTY_RESPONSE): The server closed the connection without sending any data.

original report (20:00, 20/12/2012):

Engineers are currently working on resolving a hardware fault, and are moving a number of desalo customers' sites to a new server in order to fix the issue.

While this is happening, affected sites will be down, and may show an error message when people try to access them.

Whatever error message you see, if your site is down it's likely to be due to the same server move.

Please be assured that we are working to get sites up and running again as soon as we can.

Apologies if your site is affected at the moment, and thanks for your patience.

Message from the Server Fixing Team

Due to issues with the RAID controller, we are replacing the hardware for this server and will be restoring all data from our backup server to the new hardware. Currently, we are working on connecting the backup server, and in the meantime you may be experiencing downtime with your websites or may be unable to access your server. We apologize for the inconvenience... other servers or services (such as MySQL or email) are affected.

resolved: network outage affecting some sites and email (since at least 14:35, 14/03/2011)

last updated: 14/03/2011, 16:55 (UK local time)

16:55: All sites appear to be up-and-running and functioning normally.

Please let us know if you think are still affected by this issue.

Many thanks for being so patient and understanding. Much appreciated.

15:50: Some of the sites that were down half an hour ago are back up. We're keeping a close eye on things, but really just have to wait for engineers to get a fix in place.

15:00: We're trying to get more info on this and will keep you informed on this page.

Please be assured that the host's network administrators and engineers are aware of the issue and working to fix the issue. They say they are currently operating on half a core and that partial connectivity has been restored.

We're very sorry for any outage you may be experiencing.

Please let us know if you are still affected by this issue.

Many thanks.

earlier service issues:

resolved: some sites experiencing difficulties since 11:30, 10/10/09

latest: Everything seems to be back to normal (last tested at 15:00).


Tech support staff are aware of the problem and working on it. Here is their report:

It looks like one of our edge switches is having some problems causing packetloss which is making the machines extremely slow and unresponsive. More information and possibly an ETA will come later.

It looks like some sites are already coming back online (as of 12:30), but we'll naturally keep an eye on things for you. Any concerns, please call us on 01896 820161 or mob: +447850176953. Apologies from us, and thanks for your understanding.

resolved: web and email access issue for some desalo customers on Sat 22/3/08 and Tues 25/3/08

Some of our US-hosted customers were affected by scheduled maintenance to carry out a large-scale server move and systems upgrade overnight in the US. Planned downtime was 12 hrs, and most sites were up and running again within 8 hrs, but there have been knock-on effects. Engineers are aware of these and have been working to clear up any issues.

We noticed that four of our customers' sites were down for between 20 minutes and, intermittently, around 2 hours between 20:52 and 22:45 (UK local time) on Tuesday, 25/3/08. Again, reassuringly, engineers at the data centre housing the affected cluster were aware of the fact and working on the problem before we even noticed it.

Update, 27/3/08, 09:45: Two sites were extremely sluggish for around four hours last night, but all seem to be performing optimally now, and have been since around 23:00 GMT, 26/03/08.

The issues appear to be resolved, but we'll naturally keep an even closer eye on things, just to make sure.

Update, 27/3/08, 12:01: The two sites that were sluggish last night have been fine in all of this morning's tests.

The adverse effects of last Saturday's move appear to have now been resolved. We use an external, independent monitoring service that tests the availability of sites every ten minutes, and it has not reported any problems since 25/3/08. Speed tests are also back to normal, but please let us know if you notice any issues with your own site.

web and email access issue for some desalo customers on Tue 24th July 2007 - resolved since late Tue 24th/early Wed 25th July 2007

Here's a statement from the company hosting the affected sites (who had nothing to do with the affected data centre, but were hit because the transatlantic carrier was in the centre - arrggh!):

The problem between here and the UK (and many overseas locations) seems to have stemmed from an outage with a large service provider called "Global Crossing". I believe there was a major outage with this provider, and it affected pretty much anyone who's [sic] path between themselves and our datacenters went through Global Crossing...

Global Crossing can in turn quite justifiably claim that they were not to blame, as the outage was down to backup generators not kicking in at the data centre in San Francisco that they are housed in.

Such data centres are supposed to cope with anything the world can throw at them, but this time their backups failed. Some huge sites were affected.

Slashdot had fun with the incident, especially as the date centre concerned had that very day issued a press release announcing more than 2 years without any downtime:

Again, our apologies to any of our customers who found that they were affected by the outage.

posted by Chris at 00:50

earlier announcement: 10:52, Wed 25th July, 2007

web and email access issue for some desalo customers since morning of Tue 24th July - now resolved

This issue seems to have been sorted out. The three sites that I found were affected yesterday are now working again.

Please let me know if you notice any continuing problems, and thanks again for your patience.

posted by Chris at 10:52

earlier announcement: 21:03, Tue 24th July, 2007

web and email access issue for some desalo customers since morning of Tue 24th July

This issue is still going on for some users, unfortunately.

It's not a server issue (though it still looks like a DNS problem to me, as it's consistently affecting the same domain names). On the same servers there are a) sites that are fully accessible from the UK, b) sites that are accessible from Germany or other parts of Europe, but not the UK and c) sites that are accessible from neither the UK nor other parts of Europe that I've been able to hop through from when testing.

Naturally, I'll keep watching out for news and will keep you posted.

Many thanks for your patience.

posted by Chris at 21:03

earlier announcement: 15:55, Tue 24th July, 2007

There's discussion elsewhere on the web as to whether it's a routing issue or a DNS issue, but whichever it is, the fact remains that a few people in the UK are not seeing their sites and are not able to get their mail via POP3 (i.e. they can use to access their mailboxes, but not a program such as Outlook, Thunderbird or Apple's Mail).

In contrast to our earlier announcements, it now looks as though it's not the whole of Europe that's unable to access the rather small number of affected sites, but perhaps only the UK (though this is far from clear). What is clear is that many (perhaps all) of the sites that people in the UK find it impossible to access can be accessed from other parts of Europe. There have been confirmations of such findings from Finland and Germany.

If you'd like to test whether your own site is available from Germany, try this proxy:

It's free and quick, and you don't need to register.

If you are a desalo customer and you'd like to talk any of this over, please feel free to call me in the UK on 01896 820255 or 01896 820161 at any time up to 23:00 tonight (24 July 2007) and from 10:00 tomorrow morning if things haven't been fixed by then.

We'll update you on this page as soon as things appear to have been resolved, or if we receive more information. Oh, and just to be clear, this is not something we can do anything about ourselves. Nor is it something our immediate suppliers can do anything about. All we can do at the moment is pass reports up the line, monitor the situation and report back to you here as soon as we find out more.

Many thanks for your patience.

posted by Chris at 15:55

earlier announcement: 14:52, Tue 24th July, 2007

Not all of our sites using US servers are affected. Please let us know if yours is and we'll look into it and pass reports up the line to technical support staff for them to pass on to Global Crossing. Might be due to a corrupted DNS registry as it does NOT appear to be server dependent.

We'll update you on this page as soon as things appear to have been resolved, or if we receive more information.

earlier announcement: 14:00, Tue 24th July, 2007 (edited at 14:52 to improve factual accuracy as more information became available)

Some customers are finding that their sites and emails are currently (afternoon of July 24th 2007) unavailable.

There have been reports that this is due to a transatlantic link run by a connection provider called Global Crossing ( going down and a data centre run by Dreamhost thus becoming unavailable for European visitors (though not the rest of the world).

This means that a) you may not be able to see your site from the UK at the moment, b) you may not be able to access mail from mailboxes associated with affected sites via conventional means (though may still access them via or similar services) c) that many people will still be able to see your sites, even if we in the UK can't see them just now, and will still be able to send you mail.

Dreamhost have contacted other UK customers to let them know that they are aware of the issue, but that it is, unfortunately, outside their control.

I'm not aware of any announcement as yet from Global Crossing, or if any re-routing work is underway, but can only assume that it is.

Apologies if this incident is affecting you and your sites/mail. One upside is that you should not lose any mail as a result. In our tests, we've still been able to send mail to affected domains and have been able to access mail via (or gmail for sites that forward their mail to gmail backups), just not via POP3.

That means that mail is still arriving, but that it's currently inaccessible via the usual means. That being so, it should be there waiting for you once you do get access again if you find you can't access your mail at this very moment.

We'll keep an eye on things and will let you know (here on this page) as soon as we know any more.

There are currently no other issues that we are aware of.

Please email or call us on 01896 820255 if you have specific problems to report, or would like help with support issues.


Earlier Issues, resolved since ~16:30, Wed 27th June:

Due to the severe recent flooding in Sheffield and Leeds, some customers that we provide UK web and email hosting for may have experienced technical difficulties on Wed 27th June.

Sorry if this caused you any problems.

Unfortunately, the event was beyond our control (and beyond the control of our suppliers), but we'll keep you posted on this page if we hear of further developments.

At the moment (20:33, 27/6/07) things seem to be back to normal.


Here are status reports from Donhost/Pipex, who provide the affected web and email data centres:

Current Status
These issues may have had some impact on your service recently, but they are now resolved. We apologise for any inconvenience caused.

Past Status Reports

Time 12:58 Wed 27 Jun 2007
Services Affected All servers
Details We are currently experiencing severe network connectivity issues in our Leeds datacentre, which is affecting many services for which we sincerely apologise. Engineers are attending and we are treating this with the highest urgency to get services restored as soon as possible...

...We shall update here as soon as we have any further information. Thank you for bearing with us during this time.
Update 13:28 Wed 27 Jun:
The engineers working on this issue have advised that our primary fibre link has been damaged by the severe flooding which has taken place in Sheffield over the last few days.

As of Tuesday of this week the failover link has been in use, however a line break occured at approximately 12pm Wednesday 27 June, resulting in sub-optimal routing for all Pipex customers with services terminating in our Manchester and Leeds data centres, as the third standby line is heavily loaded network traffic.

Presently engineers are working on bringing additional capacity to the final standby line while the issues relating to the Primary and Secondary are addressed.

Unfortunately, the cause of these issues are beyond the control of Donhost or Pipex, the causes for these issues namely being extreme weather conditions in this region for the Primary link, and human error caused by third party workmen on the Secondary line. We apologise for any inconvenience caused by this matter and once again assure you that our engineers are working on rectifying the problem as soon as they are able.

15:05 Wed 27 Jun:
Engineers have now routed connections to our network through back-stored 'spare' bandwidth while the issues with our primary and secondary network links are corrected. This has resulted in significant performance improvement for all of our customers. Work is ongoing to bring the network back to full capacity.

15:51 Wed 27 Jun:
Connectivity for our network has now been re-routed through spare Pipex fibre in London.

The primary network link is well on the way to being repaired in the Sheffield 'Point of Presence' and will be brought back online later this afternoon.

The location of secondary network link fault has now been identified by engineers and will be repaired fully in due course.

NEXT »: Services

Copyright ©2003 - 2021 Desalo : : Print This Page :
call us now in the UK on 01896 820255 :: Terms & Conditions